ANNUAL REPORT OF THE CENTRAL BUREAU FOR ASTRONOMICAL TELEGRAMS

1989

Activity was again at a record level during 1989, the number of Circulars issued, 240, being marginally greater than the 230 in 1987.

The high activity in 1989 was due principally to the record numbers of new comets (20) and of new Apollo/Aten objects (13) discovered. One of the latter, 1989 FC, approached the earth to within 0.0046 AU (just breaking the previous record approach—that of Hermes in 1937) about a week before its discovery. The 14 recoveries of periodic comets were topped only by the 17 of 1987; a case can be made for increasing the 1989 figure by one, for after the announcement in 1988 of the anomalous brightening of Chiron the discovery in 1989 of a tenous coma extending from this object was not a complete surprise.

Twenty-one supernovae were found to be at maximum in 1989, and additional discoveries that were at maximum in 1988 raised the number of 1988 supernova discoveries to a record 29 for one year. One new nova was discovered, and a second was immediately recognized by its discoverer as a recurrence of the 1937 nova V745 Sco; a nova discovered at x-ray wavelengths was found to be a recurrence of V404 Cyg, previously recorded at its outburst in 1938.

Announcements were made in the Circulars of the six new satellites and the ring system of Neptune discovered by Voyager 2, and the occultation of a bright star by Saturn’s satellite Titan in July was widely observed in Europe.

This was the first full year of e-mail dissemination of the Circulars, and by the year’s end there were more than 110 e-mail subscribers. This has reduced pressure on the Computer Service proper, although the frequent unreliability of e-mail continues to make this facility for remote logins to read the Circulars very desirable, and by the end of 1989 the combined e-mail/Computer Service had a total of about 180 subscribers.

The number of subscribers to the printed Circulars ranged between 706 in February and 726 in October. It is again worth noting that the printed Circulars are the “official” version, and that since the whole subscription schedule of the Bureau is predicated on the assumption that those who read the electronic Circulars will also subscribe to the printed version, the continuing decline from previous years, while understandable, is cause for concern. A note